OpenStudy (anonymous):

factor x^2+6x+9. show work

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to 9 and add to 6

jimthompson5910 (jim_thompson5910):

Those numbers are?

OpenStudy (anonymous):

set up your problem as (x+ )(x+ ) then find the numbers you can use to multiply to get 9 and then have those numbers equal to 6. once you find those numbers, you place those numbers after the plus sign in each () unit. to check you do the problem.

OpenStudy (anonymous):

use ac factoring, example: x^2+5x+6. its in the form ax^2+bx+c. so, a =1, and c= 6. thus, ac=6. we will find numbers which when we multiply we get 6, and when we add we get 5 (coefficient of the middle term).

OpenStudy (anonymous):

so, 3 and 2, since 3*2 = 6, and 3+2 = 5. then, we will use this to rewrite the middle term, thus: x^2 +2x+3x+6.

OpenStudy (anonymous):

split it into two groups, where each group has a gcf, thus, (x^2+2x)+(3x+6). factor out the gcf from each group.

OpenStudy (anonymous):

so, x(x+2)+3(x+2).factor out the common factor, so, (x+2)(x+3).